Significance level of 0.05, test whether the slope of the regression line is negative. Form a 95% confidence interval on the slope of the line.

The regression equation is, Y = 6.0 - 0.7x

predictor           coef               stdev          t-ratio

constant            6.0                2.558           2.32

X                      -0.7                0.086            -8.28

ANOVA

source           df                ss                    ms                F

regression      1               648.12                648.12           68.58

error               6                56.72                 9.45

total               7              704.84
Answer #1

Answer: We need to test whether the slope of the regression Ime is negative. n=7 dhe n-1 = 7-1=6 df = 6 This is a left talled